# Square-free orders for CM elliptic curves modulo $p$ in short intervals

### v1

@@ Abstract.
Let $E$ be a CM elliptic curve over $\mathbb{Q}$. We refine the work of Cojocaru on the asymptotic formulae for the number of primes $p\leq x$ for which the reduction modulo $p$ of $E$ is of square-free order. Also, we derive an unconditional short interval variant for the asymptotics. Compared to the estimate derived from the generalised Riemann hypothesis, the presented result is valid for even shorter intervals. Furthermore, we improve the short interval variant of the cyclicity problem for CM elliptic curves obtained by the author.
@@
Statements and Declarations: The author was an NCTS postdoctoral fellow, and he was also supported by a PIMS postdoctoral fellowship and the University of Lethbridge.
@@ 1. Introduction
Let $E$ be an elliptic curve defined over $\mathbb{Q}$, and let $N_{E}$ be its conductor. For a prime $p$ of good reduction, we let $\bar{E}$ be the reduction of $E$ modulo $p$ and $\bar{E}(\mathbb{F}_{p})$ be the group of rational points of $\bar{E}$ over $\mathbb{F}_{p}$. By Hasse’s bound, there is an integer $a_{p}$ such that $|a_{p}|\leq 2\sqrt{p}$ and $|\bar{E}(\mathbb{F}_{p})|=p+1-a_{p}$. For $p\geq 5$, $p$ is called supersingular if $a_{p}=0$; otherwise, $p$ is ordinary.
@@
[…]
@@
where
@@
Also, by [5, Theorem 1.3], one knows $\delta_{E}>0$ whenever $K$ is $\mathbb{Q}(\sqrt{-11})$, $\mathbb{Q}(\sqrt{-19})$, $\mathbb{Q}(\sqrt{-43})$, $\mathbb{Q}(\sqrt{-67})$, or $\mathbb{Q}(\sqrt{-163})$.
@@
[…]
@@
one may ask for an asymptotic formula of the distribution of primes in short intervals when $(x,2x]$ is replaced by $(x,x+h]$ for some $x^{1-\delta}\leq h\leq x$ with $\delta\in[0,1)$. Under the Riemann hypothesis, (8) leads to that for any $A>0$,
@@
whenever $x^{1-\delta}\leq h\leq x$ with $\delta\in[0,\frac{1}{2})$. Although the Riemann hypothesis is still out of reach, there are several progresses towards (9). For example, Huxley [11] showed that (9) is valid for $x^{1-\delta}\leq h\leq x$ with $\delta\in[0,\frac{5}{12})$. From the above discussion, one may further consider
[…]
@@
The proofs of (2) and (4) in [5] respectively rely on the effective versions of the Chebotarev density theorem established by Lagarias-Odlyzko [13] and M.R. Murty-V.K. Murty-Saradha [16]. In contrast, our proofs of (5) and Theorem 2 make a use of the Bombieri-Vinogradov theorem for number fields due to Huxley [10] (see also Theorem 4 below) and its short interval variant (Theorem 5). This is inspired by the works of Akbary-V.K. Murty [1] and M.R. Murty [15] on the cyclicity problem discussed below. The key observation is that the Chebotarev conditions involved in the proofs can be translated into the congruence conditions over certain ray class groups via Artin reciprocity. Nonetheless, to establish (6), we still need to require the effective result of M.R. Murty, V.K. Murty, and Saradha [16]. The new input in our argument is to consider a finer splitting (17) below and invoke the Brun-Titchmarsh inequality due to Hinz and Lodemann [9] to control the “middle range” $\mathcal{N}_{0}(x,y,x^{\theta})$.
@@
(12) $\pi_{c}(h+h,E)-\pi_{c}(x,E)=\mathfrak{c}_{E}(\operatorname{Li}(x+h)-\operatorname{Li}(x))+O_{E}\Big{(}\frac{h}{(\log x)^{A}}\Big{)}$
for any $x^{1-\delta}\leq h\leq x$ with $0\leq\delta<\frac{1}{4}$. Unconditionally, in [18], the author showed that the estimate (12) is valid whenever $x^{1-\delta}\leq h\leq x$, with $0\leq\delta<\frac{1}{25}$. In this article, we shall further sharpen such an unconditional result as follows.
@@
The rest of this article is arranged as follows. In the next section, we will discuss the strategy of proving Theorems 1 and 2. We will prove Theorems 1, 2, and 3 in Sections 3, 4, 5, respectively. We also note that throughout our poof of (5) and Theorems 2 and 3, we will implicitly assume $N_{E}\leq\exp(\sqrt{\log x})$ as the claimed estimates hold trivially otherwise.
2. Overview of the proofs of Theorems 3 and 4
@@
for any prime $\mathfrak{p}\nmid mN_{E}$ of $K$, where $\sigma_{\mathfrak{p}}$ denotes the Artin symbol at $\mathfrak{p}$, ${\rm N}(\mathfrak{p})$ denotes the norm of $\mathfrak{p}$, and $a_{\mathfrak{p}}={\rm N}(\mathfrak{p})+1-|\bar{E}(\mathbb{F}_{\mathfrak{p}})|$. Following Cojocaru [5], we set
@@
By a criterion of Deuring [7], for any prime $p\geq 5$ that is of good reduction for $E$, $p$ is ordinary (i.e. $a_{p}\neq 0$) if and only if $p$ splits completely in $K$. Moreover, if $p$ splits completely in $K$ and $\mathfrak{p}$ is a prime of $K$ above $p$, then we have ${\rm N}(\mathfrak{p})=p$ and $a_{\mathfrak{p}}=a_{p}$. Thus, from (13) and (14), it follows that the condition $|\bar{E}(\mathbb{F}_{p})|\equiv 0(\operatorname{mod}m^{2})$ is equivalent to $\sigma_{\mathfrak{p}}\subseteq D_{m^{2}}.$ Hence, we obtain
#$\#\{p\leq x\mid p\nmid mN_{E},a_{p}\neq 0,|\bar{E}(\mathbb{F}_{p})|\equiv 0(\operatorname{mod}m^{2})\}$
@@
where the error comes from the degree-two primes $\mathfrak{p}$ of $K$ such that ${\rm N}(\mathfrak{p})\leq x$ (there are at most $O(\sqrt{x}/\log x)$ of such primes) and ramified primes (cf. [1, Eq. (3.2)]).
@@
(15) $\pi_{D_{m^{2}}}(x)=\#\{\mathfrak{p}\subseteq\mathcal{O}_{K}\mid{\rm N}(\mathfrak{p})\leq x,\mathfrak{p}\nmid mN_{E},\sigma_{\mathfrak{p}}\subseteq D_{m^{2}}\}.$
@@
where $h_{E}(x,\mathbb{Q})$ is defined in (1),
@@
To establish asymptotic formulae for $h_{E}(x,\mathbb{Q})$ and $h_{E}(x,h,\mathbb{Q})$, it is crucial to analyse $\mathcal{N}(x,y)$ and $\mathcal{N}(x,h,y)$, respectively. In [5], Cojocaru applied effective versions of the Chebotarev density theorem established in [13, 16] to handle $h_{E}(x,\mathbb{Q})$. Under GRH, the effective version due to M.R. Murty-V.K. Murty-Saradha [16] provides a satisfactory error term. In contrast, the unconditional result obtained by Lagarias and Odlyzko gives a much weaker estimate, which led Cojocaru to use a more delicate sieve argument. In light of the work of Akbary-V.K. Murty [1] and M.R. Murty [15], we observe that the Chebotarev conditions involved (i.e., $\sigma_{\mathfrak{p}}\subseteq D_{m^{2}}$) can be further translated into the congruence conditions over certain ray class groups via Artin reciprocity (see Section 3 for more details). This observation hints at using the following Bombieri-Vinogradov theorem for number fields and its short interval variant. As shall be seen, such a Bombieri-Vinogradov theorem for number fields will play a crucial role in improving Cojocaru’s estimate (2).
@@
where $\mathfrak{p}\sim\mathfrak{a}\enspace(\operatorname{mod}\mathfrak{q})$ means that $\mathfrak{p}$ and $\mathfrak{a}$ belong to the same ray class of the ray class group modulo $\mathfrak{q}$. (For the background on ray class groups, we refer the reader to [3, Ch. 3].) In [10], Huxley proved the following Bombieri-Vinogradov theorem for number fields, which improves the previous result of Wilson [17].
@@
Also, we recall the following short interval variant established by the author [18].
@@
In our consideration, $F$ will be an imaginary quadratic field of class number one, and the factor $\frac{h(\mathfrak{q})}{\phi(\mathfrak{q})}$ can be dropped from the estimates in Theorems 4 and 5 for this case as follows. Let $h_{F}$ denote the class number of $F$ and $r_{1}$ be the number of real embeddings of $F$. It is known that
@@
where $T(\mathfrak{q})$ is the number of residue classes $(\operatorname{mod}\mathfrak{q})$ that contain a unit (see [17, Eq. (7)]). Note that for imaginary quadratic fields, $T(\mathfrak{q})\leq 6$ and hence $\frac{h(\mathfrak{q})}{\phi(\mathfrak{q})}\geq\frac{1}{6}$. Thus, the factor $\frac{h(\mathfrak{q})}{\phi(\mathfrak{q})}$ can be removed.
@@ 3.1. Proof of the first part of Theorem 1
Now, we let $\tilde{D}_{m^{2}}$ be the conjugacy set in $\operatorname{Gal}(K_{\mathfrak{fm^{2}}}/K)$ obtained from $D_{m^{2}}$ through the quotient map from $\operatorname{Gal}(K_{\mathfrak{fm^{2}}}/K)$ to $\operatorname{Gal}(K(E[m^{2}])/K)$, and we let $\pi_{\tilde{D}_{m^{2}}}(x)$ denote the number of unramified primes $\mathfrak{p}$ in $K_{\mathfrak{fm^{2}}}/K$ such that ${\rm N}(\mathfrak{p})\leq x$ and the Artin symbol $\tilde{\sigma}_{\mathfrak{p}}$ at $\mathfrak{p}$ is contained in $\tilde{D}_{m^{2}}$. It is known that
@@
as $n_{K}=2$ and $[K_{\mathfrak{fm^{2}}}:K]\leq h(\mathfrak{fm^{2}})\leq 1\cdot 2^{0}\cdot\phi(\mathfrak{fm^{2}})\leq{\rm N}(\mathfrak{fm^{2}})$. Hence, we arrive at
@@
Thus, by Theorem 4, together with the discussion below Theorem 5, for $z^{4}\leq x^{1/2-\epsilon}$ and $A>0$, we have
@@ 3.2. Proof of the second part of Theorem 1
In [5, Lemma 3.12], Cojocaru showed that $|D_{q^{2}}|\leq q^{2}$ for any odd prime $q$ that is unramified in $K$. From which, she deduced in [5, Corollary 3.13] that $|D_{m}^{2}|\leq m^{2}$ for any odd positive square-free integer $m$ composed of primes which are unramified in $K$. We remark that this bound can be extended to all square-free integers $m>1$ so that $|D_{m^{2}}|\ll m^{2}$ as follows.iiiiiiIn fact, such an extension has been used implicitly in [5, Eq. (34) and (39)]. Observe that for any prime $\ell$ that is even or ramifies in $K$, we always have
@@
As only finitely many primes ramify in $K$, we can write this bound as $|D_{\ell^{2}}|\ll_{K}1$. In addition, since $K$ must be one of night imaginary quadratic fields of class number one, we deduce that $|D_{\ell^{2}}|$ is absolutely bounded (i.e. $|D_{\ell^{2}}|\ll 1$) for any prime $\ell$ that is even or ramifies in $K$. This, combined with the above-mentioned result of Cojocaru, yields $|D_{m}^{2}|\ll m^{2}$ for any square-free integer $m>1$. As a direct consequence, we have
@@
where we used the bound $[K(E[k]):K]\gg\phi(k)^{2}$ for any integer $k\geq 3$ (see, e.g., [5, Eq. (19)]). Hence, the last sum in (27) is $\ll\frac{x}{y\log x}$ (cf. [5, Eq. (39)]).

### v2

@@ Abstract.
Let $E$ be a CM elliptic curve over $\mathbb{Q}$. We refine the work of Cojocaru on the asymptotic formulae for the number of primes $p\leq x$ for which the reduction modulo $p$ of $E$ is of square-free order. Also, we derive an unconditional short interval variant for the asymptotics. Compared to the estimate derived from the generalised Riemann hypothesis, the presented result is valid for even shorter intervals. Furthermore, we improve the short interval variant of the cyclicity problem for CM elliptic curves previously obtained by the author.
@@
Statements and Declarations: The author was an NCTS postdoctoral fellow, and he was also supported by a PIMS postdoctoral fellowship and the University of Lethbridge during conducting part of this research. He is currently supported by the NSTC grant 114-2628-M-110-006-MY4.
@@ 1. Introduction
Let $E$ be an elliptic curve defined over $\mathbb{Q}$, and let $N_{E}$ be its conductor. For a prime $p$ of good reduction, we let $\bar{E}$ be the reduction of $E$ modulo $p$ and $\bar{E}(\mathbb{F}_{p})$ be the group of rational points of $\bar{E}$ over $\mathbb{F}_{p}$. By Hasse’s bound, there is an integer $a_{p}$ such that $|a_{p}|\leq 2\sqrt{p}$ and $|\bar{E}(\mathbb{F}_{p})|=p+1-a_{p}$. For $\geq 5, pp$ is called supersingular if $a_{p}=0$; otherwise, $p$ is ordinary.
@@
(1) $h_{E}(x,\mathbb{Q})=\#\{p\leq x\mid\text{$p\nmid N_{E}$, $a_{p}\neq 0$, and $|\bar{E}(\mathbb{F}_{p})|$ is square-free}\}$
@@
where $\operatorname{Li}(x)=\int_{2}^{x}\frac{dt}{\log t}$ is the logarithmic integral function,
@@
Recall that as $E$ is defined over $\mathbb{Q}$, the field $K$ is one of the nine imaginary quadratic fields of class number 1. It worth remarking that by [5, Theorem 1.3], one knows $\delta_{E}>0$ whenever $K$ is $\mathbb{Q}(\sqrt{-11})$, $\mathbb{Q}(\sqrt{-19})$, $\mathbb{Q}(\sqrt{-43})$, $\mathbb{Q}(\sqrt{-67})$, or $\mathbb{Q}(\sqrt{-163})$.
@@
$\pi_{E}^{\mathrm{SF}}(x)=\#\{p\leq x\mid\text{$p\nmid N_{E}$ and $|\bar{E}(\mathbb{F}_{p})|$ is square-free}\}.$
@@
one may ask for an asymptotic formula of the distribution of primes in short intervals when $(x,2x]$ is replaced by $(x,x+h]$ for some $x^{1-\delta}\leq h\leq x$ with $\delta\in[0,1)$. Under the Riemann hypothesis, (8) implies that for any $A>0$,
@@
whenever $x^{1-\delta}\leq h\leq x$ with $\delta\in[0,\frac{1}{2})$. Although the Riemann hypothesis is still out of reach, there are some progresses towards (9). For example, Huxley [11] showed that (9) is valid for $x^{1-\delta}\leq h\leq x$ with $\delta\in[0,\frac{5}{12})$. From the above discussion, one may further consider
(10) $h_{E}(x,h,\mathbb{Q})=\#\{x<p\leq x+h\mid\text{$p\nmid N_{E}$, $a_{p}\neq 0$, and $|\bar{E}(\mathbb{F}_{p})|$ is square-free}\}.$
@@
The proofs of (2) and (4) in [5] respectively rely on the effective versions of the Chebotarev density theorem established by Lagarias-Odlyzko [13] and M.R. Murty-V.K. Murty-Saradha [16]. In contrast, our proofs of (5) and Theorem 2 make a use of the Bombieri-Vinogradov theorem for number fields due to Huxley [10] (see also Theorem 4 below) and its short interval variant (Theorem 5). This is inspired by the works of Akbary-V.K. Murty [1] and M.R. Murty [15] on the cyclicity problem discussed below. The key observation is that the Chebotarev conditions involved in the proofs can be translated into congruence conditions over certain ray class groups via Artin reciprocity. Nonetheless, to establish (6), we still require the effective result of M.R. Murty, V.K. Murty, and Saradha [16]. The new input in our argument is to consider a finer splitting (17) below and invoke the Brun-Titchmarsh inequality (for $K$) due to Hinz and Lodemann [9] to control the “middle range” $\mathcal{N}_{0}(x,y,x^{\theta})$.
@@
(12) $\pi_{c}(x+h,E)-\pi_{c}(x,E)=\mathfrak{c}_{E}(\operatorname{Li}(x+h)-\operatorname{Li}(x))+O_{E}\Big(\frac{h}{(\log x)^{A}}\Big)$
for any $x^{1-\delta}\leq h\leq x$ with $0\leq\delta<\frac{1}{4}$. Unconditionally, in [18], the present author showed that the estimate (12) is valid whenever $x^{1-\delta}\leq h\leq x$, with $0\leq\delta<\frac{1}{25}$. In this article, we shall further sharpen such an unconditional result as follows.
@@
The rest of this article is arranged as follows. In the next section, we will discuss the strategy of proving Theorems 1 and 2. We will prove Theorems 1, 2, and 3 in Sections 3, 4, 5, respectively. We also note that throughout our proofs of (5) and Theorems 2 and 3, we will implicitly assume $N_{E}\leq\exp(\sqrt{\log x})$ as the claimed estimates hold trivially otherwise.
2. Overview of the proofs of Theorems 1 and 2
@@
for any prime $\mathfrak{p}\nmid mN_{E}$ of $K$, where $\sigma_{\mathfrak{p}}$ denotes the Artin symbol at $\mathfrak{p}$, ${\rm N}(\mathfrak{p})$ is the norm of $\mathfrak{p}$, and $a_{\mathfrak{p}}={\rm N}(\mathfrak{p})+1-|\bar{E}(\mathbb{F}_{\mathfrak{p}})|$. Following Cojocaru [5], we set
@@
By a criterion of Deuring [7], for any prime $p\geq 5$ that is of good reduction for $E$, $p$ is ordinary (i.e. $a_{p}\neq 0$) if and only if $p$ splits in $K$. Moreover, if $p$ splits in $K$ and $\mathfrak{p}$ is a prime of $K$ above $p$, then we have ${\rm N}(\mathfrak{p})=p$ and $a_{\mathfrak{p}}=a_{p}$. Thus, from (13) and (14), it follows that the condition $|\bar{E}(\mathbb{F}_{p})|\equiv 0(\operatorname{mod}m^{2})$ is equivalent to $\sigma_{\mathfrak{p}}\subseteq D_{m^{2}}.$ Hence, we obtain
$\#\{p\leq x\mid p\nmid mN_{E},a_{p}\neq 0,|\bar{E}(\mathbb{F}_{p})|\equiv 0(\operatorname{mod}m^{2})\}$
@@
where the error comes from the degree-two primes $\mathfrak{p}$ of $K$ such that ${\rm N}(\mathfrak{p})\leq x$ (there are at most $O(\sqrt{x}/\log x)$ such primes) and ramified primes (cf. [1, Eq. (3.2)]).
@@
(15) $\pi_{D_{m^{2}}}(x)=\#\{\mathfrak{p}\subset\mathcal{O}_{K}\mid{\rm N}(\mathfrak{p})\leq x,\mathfrak{p}\nmid mN_{E},\sigma_{\mathfrak{p}}\subseteq D_{m^{2}}\}.$
@@
where $h_{E}(x,\mathbb{Q})$ is as introduced in (1),
@@
To establish asymptotic formulae for $h_{E}(x,\mathbb{Q})$ and $h_{E}(x,h,\mathbb{Q})$, it is crucial to analyse $\mathcal{N}(x,y)$ and $\mathcal{N}(x,h,y)$, respectively. In [5], Cojocaru applied effective versions of the Chebotarev density theorem established in [13, 16] to handle $h_{E}(x,\mathbb{Q})$. Under GRH, the effective version due to M.R. Murty-V.K. Murty-Saradha [16] provides a satisfactory error term. In contrast, the unconditional result obtained by Lagarias and Odlyzko gives a much weaker estimate, which led Cojocaru to use a more delicate sieve argument. In light of the works of Akbary-V.K. Murty [1] and M.R. Murty [15], we observe that the Chebotarev conditions involved (i.e., $\sigma_{\mathfrak{p}}\subseteq D_{m^{2}}$) can be further translated into congruence conditions over certain ray class groups via Artin reciprocity (see Section 3 for more details). This observation hints at using the following Bombieri-Vinogradov theorem for number fields and its short interval variant. As shall be seen, such a Bombieri-Vinogradov theorem for number fields will play a crucial role in improving Cojocaru’s estimate (2).
@@
where $\mathfrak{p}\sim\mathfrak{a}\kern 5.0pt(\operatorname{mod}\mathfrak{q})$ means that $\mathfrak{p}$ and $\mathfrak{a}$ belong to the same ray class of the ray class group modulo $\mathfrak{q}$. (For the background on ray class groups, we refer the reader to [3, Ch. 3].) In [10], Huxley proved the following Bombieri-Vinogradov theorem for number fields, which improves the previous result of Wilson [17].
@@
Also, we recall the following short interval variant established by the present author in [18].
@@
In our consideration, we will take $F=K$, an imaginary quadratic field of class number one associated to a CM elliptic curve $E$ as discussed in the introduction. For such an instance, the factor $\frac{h(\mathfrak{q})}{\phi(\mathfrak{q})}$ can be dropped from the estimates in Theorems 4 and 5 as follows. Let $h_{F}$ denote the class number of $F$ and $r_{1}$ be the number of real embeddings of $F$. It is known that
@@
where $T(\mathfrak{q})$ is the number of residue classes $(\operatorname{mod}\mathfrak{q})$ that contain a unit (see [17, Eq. (7)]). Now, for imaginary quadratic fields of class number one, it is known that $T(\mathfrak{q})\leq 6$, and hence $\frac{h(\mathfrak{q})}{\phi(\mathfrak{q})}\geq\frac{1}{6}$. Thus, the factor $\frac{h(\mathfrak{q})}{\phi(\mathfrak{q})}$ can be removed.
@@ 3.1. Proof of the first part of Theorem 1
Now, we let $\tilde{D}_{m^{2}}$ stand for the preimage of $D_{m^{2}}$ under the quotient map from $\operatorname{Gal}(K_{\mathfrak{fm^{2}}}/K)$ to $\operatorname{Gal}(K(E[m^{2}])/K)$. As $\tilde{D}_{m^{2}}$ is a conjugacy set in $\operatorname{Gal}(K_{\mathfrak{fm^{2}}}/K)$, we can let $\pi_{\tilde{D}_{m^{2}}}(x)$ denote the number of unramified primes $\mathfrak{p}$ in $K_{\mathfrak{fm^{2}}}/K$ such that ${\rm N}(\mathfrak{p})\leq x$ and the Artin symbol $\tilde{\sigma}_{\mathfrak{p}}$ at $\mathfrak{p}$ is contained in $\tilde{D}_{m^{2}}$. It is known that
@@
as $n_{K}=2$ and $K_{\mathfrak{fm^{2}}}:K]\leq h(\mathfrak{fm^{2}})\leq 1\cdot 2^{0}\cdot\phi(\mathfrak{fm^{2}})\leq{\rm N}(\mathfrak{fm^{2}})$. Hence, we arrive at
@@
Therefore, by Theorem 4, together with the discussion beneath Theorem 5, for $z^{4}\leq x^{1/2-\epsilon}$ and $A>0$, we have
@@ 3.2. Proof of the second part of Theorem 1
In [5, Lemma 3.12], Cojocaru showed that $|D_{q^{2}}|\leq q^{2}$ for any odd prime $q$ that is unramified in $K$. From this, she deduced in [5, Corollary 3.13] that $|D_{m^{2}}|\leq m^{2}$ for any odd positive square-free integer $m$ composed of primes which are unramified in $K$. We remark that this bound can be extended to all square-free integers $m>1$ so that $|D_{m^{2}}|\ll m^{2}$ as follows.iiiiiiIn fact, such an extension has been used implicitly in [5, Eq. (34) and (39)]. Observe that for any prime $\ell$ that is even or ramifies in $K$, we always have
@@
As only finitely many primes ramify in $K$, we can write this bound as $|D_{\ell^{2}}|\ll_{K}1$. In addition, since $K$ must be one of nine imaginary quadratic fields of class number one, we deduce that $|D_{\ell^{2}}|$ is absolutely bounded (i.e. $|D_{\ell^{2}}|\ll 1$) for any prime $\ell$ that is even or ramifies in $K$. This, combined with the above-mentioned result of Cojocaru, yields $|D_{m^{2}}|\ll m^{2}$ for any square-free integer $m>1$. As a direct consequence, we have
@@
where we used the bound $K(E[k]):K]\gg\phi(k)^{2}$ for any integer $k\geq 3$ (see, e.g., [5, Eq. (19)]). Hence, the last sum in (27) is $\ll\frac{x}{y\log x}$ (cf. [5, Eq. (39)]).